a chmist mixes 100ml of a 28% solution of acohol with another sample of 40% alcohol solution to obtain a sample containing 36% alcohol. how much of the 40% alcohol was used
Answer by mananth(16946)

------ percent ---------------- quantity
Alcohol I 28 ----------------100 ml
Alcohol II 40 ----------------x ml
Total 36 ---------------- 100 + x ml
28*100+40*x=36(100+x)
2800+40 x=3600+36x
40x-36x =3600-2800
4x=800
/ 4 x = 200 ml 40 % Alcohol II